OpenStudy (erinweeks):

PLEASEEE HELPP ! Simplify the rational expression. State any excluded values (x^2-36)/(6-x)

OpenStudy (anonymous):

Use the following property\[ a^2 -b^2 = (a+b)(a-b)\]So we have, \[ x^2 -36 = (x+6)(x-6)\] Plug this into the expression and see what cancels out

OpenStudy (erinweeks):

\[\frac{ x ^{2}- 36 }{ 6-x }\] thats the expression

OpenStudy (anonymous):

hmm couldve sworn i did this problem a few min ago lol

OpenStudy (erinweeks):

Answers A. \[-x-6 ; x \neq 6\] B. \[x + 6 ; x \neq -6\] C. \[\frac{ 1 }{ x-6 } ; x \neq 6\]

OpenStudy (anonymous):

after you plug in \[x^2-36 = (x+6)(x-6)\]you'd get\[\frac{(x+6)(x-6)}{(6-x)}\]Since\[(6-x) = -1 \times(x-6)\]so after cancelling out you'd get\[-(x+6)\]

OpenStudy (erinweeks):

okay so what answer would that be cause i dont see it ?

OpenStudy (anonymous):

option A) \[-(x+6)\] or \[-x-6\]

OpenStudy (erinweeks):

okay thank you